OpenStudy (anonymous):

A cannery processed 605 pounds of strawberries in 3.5 hours. The cannnery processed 2100 pounds in 10 hours. Write a linear equation to model the weight of strawberries "S" processed in "T" hours. How many pounds of strawberries can be processed in 12 hours?

OpenStudy (anonymous):

points are (T,S) = (time, pounds), so you have (3.5, 605) and (10, 2100). Slope is (2100-605) / (10-3.5) = 1495/6.5 = 230. T is the independent variable (like x normally is in linear equations), so that narrows it down to choice A or choice D due to the slope. You can solve for b in y=mx+b or S = mT + b by plugging in one of the points: 2100 = 230(10) + b ===> b = -200. Therefore S = 230T - 200; S(12) = 230*12 - 200 = 2560 lbs [B